# Multiverse: Transactional Memory with Dynamic Multiversioning

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: ; ; $c$s: Computing methodologies Concurrent algorithms
@@ 4. Implementation Details
6 while vNode.tbd and vNode.timestamp < rclock
@@ 4.1. Mode Q Code Paths
Versioned readers in local Mode Q (modeQ_versionedRead in LABEL:alg:tm-read) begin by checking if the address is versioned. If the address is versioned then the transaction will traverse the version list beginning from the newest version to find a suitable version with a timestamp less than or equal to its versioned timestamp. This traversal is blocked if the newest version is suitable but marked TBD. If a suitable version is found the data is returned otherwise the transaction is aborted.
@@
29modeU_versionedRead(addr):
@@
38 if lockState.version == lastVer and data == lastVal
39 return lastVal
@@
41 lastVal = *addr
42 goto retry
43 data = *addr
44 lockState2 = reread lock
45 if lockState2.version != lockState.version
46 abort()
47 return data
Algorithm 4: Pseudocode for TM Read
@@ 4.2. Mode U Code Paths
When a read-only transaction is in local Mode U (modeU_versionedRead in LABEL:alg:tm-read), it still begins as unversioned and the execution of unversioned reads is the same as in Mode Q. When reading an address, versioned read-only transactions in local Mode U still need to check if the address is versioned. If the address is already versioned then we perform the same traversal of the version list as in Mode Q.
If a versioned transaction in Mode U encounters an unversioned address then we know that there has not been any writes to this address since the transition to Mode U, otherwise, some other writer would have already versioned the address. However, we check if an address is versioned before reading the data. Since these steps are not done atomically together, it is possible for a concurrent transaction to update the address after we observe that it is unversioned, which could result in the versioned transaction observing inconsistent data. To prevent this, if a versioned transaction encounters an unversioned address it will read and make a local copy of the associated lock state. If it is unlocked, the versioned transaction must read the data at the address then reread the lock. If the lock version has not changed then we know that no concurrent updates to this address have occurred so it is safe to return the data that we read. If lock version has changed then the transaction must abort.
In the case that the (unversioned) address is locked then the versioned transaction must read and makes a local copy of the data then re-checks if the address is versioned. If the address is still unversioned, the transaction must redo the reads of the lock and data. An address in Mode U can be locked iff a writer is concurrently updating the address or due to a lock table collision. If the address is still unversioned but we observe a change in the lock version, then the address must have been locked as a result of a lock table collision, since otherwise, the writer holding the lock would have versioned the address. Alternatively, if the address is still unversioned and we do not observe a change in the lock version or the data, then the lock could be held by a transaction seeking to update this address. However, in this case, our first read of the data must have occurred before any such update, since again, otherwise, the address would have been versioned.

### v4

@@ Abstract.
journalyear: 2026copyright: ccconference: ; ; $c$s: Computing methodologies Concurrent algorithms11footnotetext: This version of the paper corrects erratum in Section 4.1 regarding version list traversals and in Section 4.2 regarding the optimization when Mode U reads encounter an unversioned address. This version of the paper also provides an expanded description of unversioning addresses in Section 4.4.
@@ 4. Implementation Details
6 while vNode.tbd and vNode.timestamp <= rclock
@@ 4.1. Mode Q Code Paths
Versioned readers in local Mode Q (modeQ_versionedRead in LABEL:alg:tm-read) begin by checking if the address is versioned. If the address is versioned then the transaction will traverse the version list beginning from the newest version to find a suitable version with a timestamp less its versioned timestamp. This traversal is blocked if the newest version is suitable but marked TBD. If a suitable version is found the data is returned otherwise the transaction is aborted.
@@
Algorithm 4: Pseudocode for TM Read and ModeQ TM Read
1modeU_versionedRead(addr):
2 lastVer = -1, lastVal = null, didRetry = false
3 retry:
@@
10 validVer = lockState.version < rClock or
11 (firstObsModeUTs != -1 and
12 firstMode2Ts < rClock)
13 if didRetry
14 verChanged = lockState.version != lastVer
15 valChanged = val != lastVal
16 if verChanged return lastVal
17 if lockState.locked and validVer and not verChanged and not valChanged
18 return lastVal
19 if lockState.unlocked and validVer return lastVal
20 abort()
21 else
@@
24 lastVal = val
25 didRetry = true
26 else if validVer return val
27 else abort()
Algorithm 5: Pseudocode for ModeU TM Read
@@ 4.2. Mode U Code Paths
When a read-only transaction is in local Mode U (modeU_versionedRead in LABEL:alg:tm-read2), it still begins as unversioned and the execution of unversioned reads is the same as in Mode Q. When reading an address, versioned read-only transactions in local Mode U still need to check if the address is versioned. If the address is already versioned then we perform the same traversal of the version list as in Mode Q.
If a versioned transaction in Mode U encounters an unversioned address then we know that there has not been any writes to this address since the transition to Mode U, otherwise, some other writer would have already versioned the address. However, we check if an address is versioned before reading the data. Since these steps are not done atomically together, it is possible for a concurrent transaction to update the address after we observe that it is unversioned, which could result in the versioned transaction observing inconsistent data. To prevent this, if a versioned transaction encounters an unversioned address it will read and make a local copy of the data at the address and the associated lock state. If the address is unlocked, then it is safe to return the data that we read so long when the address becomes versioned its initial version will have a timestamp that is less than the transactions versioned timestamp. This is the case if the lock version is less than the transaction’s versioned timestamp or if the the first observed Mode U timestamp is valid and is less than the transaction’s versioned timestamp.
In the case that the (unversioned) address is locked then the versioned transaction re-checks if the address is versioned. If the address is still unversioned, the transaction must redo the reads of the lock and data. An address in Mode U can be locked iff a writer is concurrently updating the address or due to a lock table collision. If the address is still unversioned but we observe a change in the lock version, then the address must have been locked as a result of a lock table collision, since otherwise, the writer holding the lock would have versioned the address. Alternatively, if the address is still unversioned and we do not observe a change in the lock version or the data, then the lock could be held by a transaction seeking to update this address. However, in this case, our first read of the data must have occurred before any such update, since again, otherwise, the address would have been versioned. In this case it is safe to return the first value that the transaction read so long as when the address is versioned it will be versioned at a timestamp that is less than the transactions versioned timestamp. The requirements for this to be the case are the same as in the case when the address as unlocked.
@@ 4.4. How to Unversion
There is a subtle scenario in which a single address should be unversioned (instead of unversioning the entire VLT bucket). More specifically, the version list of a single address must be invalidated or removed if the address is reclaimed while it is still versioned. This is necessary since otherwise, the address could be returned from a future allocation and would appear to still be versioned but the associated version list could contain versions that point to freed memory.